gfggggg:

What is the value of n in the equation –negative StartFraction one-half EndFraction left-parenthesis 2 n plus 4 right-parenthesis plus 6 equals negative 9 plus 4 left parenthesis 2 n plus 1 right-parenthesis.(2n + 4) + 6 = –9 + 4(2n + 1)?

ihy:

The value of n is 1

Bot1:

the value of 𝑛 n is 1.

Bot1:

damn

adrianluvvsyouu2:

(2n + 4) + 6 = –9 + 4(2n + 1) Lets subtract 6 from both sides: 2n + 4 = 4(2n + 1) - 15 Distributive property on 4(2n + 1): 2n + 4 = 8n + 4 - 15 Cancel the 4 on both sides (subtract 4): 2n = 8n - 15 Add 15 to both sides: 2n + 15 = 8n Subtract 2n from both sides: 15 = 6n Divide both sides by 6: n = 15/6 which can be simplified to 5/2 or 2.5 Whoever is saying n=1 probably googled it and got it incorrect.

skitzzz:

mathness

NaiNoah:

@skitzzz wrote:
mathness
the value of n is not 1

surjithayer:

\[-\frac{ 1 }{ 2}\left( 2n+4 \right)+6=-9+4\left( 2n+1 \right)\] multiply by 2 -(2n+4)+12=-18+8(2n+1) -2n-4+12=-18+16n+8 -2n-16n=-18+8+4-12 -18n=-18 \[n=\frac{ -18 }{ -18 }=1\] n=1

ykbatman:

n= 1

MysticEcho:

it is n=1

grimreaper124:

what they said